# Sketch-Guided Constrained Decoding for Boosting Blackbox Large Language Models without Logit Access

### v1

@@ Abstract
Constrained decoding, a technique for enforcing constraints on language model outputs, offers a way to control text generation without retraining or architectural modifications. Its application is, however, typically restricted to models that give users access to next-token distributions (usually via softmax logits), which poses a limitation with blackbox large language models (LLMs). This paper introduces sketch-guided constrained decoding (SGCD), a novel approach to constrained decoding for blackbox LLMs, which operates without access to the logits of the blackbox LLM. SGCD utilizes a locally hosted auxiliary model to refine the output of an unconstrained blackbox LLM, effectively treating this initial output as a “sketch” for further elaboration. This approach is complementary to traditional logit-based techniques and enables the application of constrained decoding in settings where full model transparency is unavailable. We demonstrate the efficacy of SGCD through experiments in closed information extraction and constituency parsing, showing how it enhances the utility and flexibility of blackbox LLMs for complex NLP tasks. [^1] We will release the code and data upon publication.
@@ 1 Introduction
Figure 1: Overview of sketch-guided constrained decoding (SGCD). In the initial sketching phase, a blackbox LLM generates a preliminary “sketch” answer without applying any constraints. Then, in the constrained generation phase, an auxiliary model, the constrained generator, refines the sketch. The refined, final output respects the specified constraints by construction.
@@
Contributions. In response, we present sketch-guided constrained decoding (SGCD), which bypasses the need for direct logit access. SGCD uses a locally hosted auxiliary model to refine the outputs of a blackbox LLM while respecting the specified constraints. We validate our method on closed information extraction and constituency parsing, two challenging NLP tasks that require the model to generate highly structured and grounded outputs. Our experiments show that SGCD significantly boosts the performance of LLMs and beats previous approaches by a wide margin.
@@ 2 Method
SGCD splits the constrained decoding task into two distinct phases: sketching and constrained generation.
@@
Constrained generation is done by a constrained generator, a smaller-scale, locally hosted LLM $P_{\text{cg}}$. Given an instruction $I_{\text{cg}}$, a set of input–sketch–output demonstrations $D_{\text{cg}}=\{(x^{i},y^{i},z^{i})\}_{i=1}^{n}$, the original input $x$, and the sketch $y^{*}$, it refines $y^{*}$ into
@@
The sketcher’s output $y^{*}$ is typically of high quality, encapsulating the necessary information for the constrained generator to produce the final sequence $z^{*}$ that adheres to the constraints $C$. Given the quality of $y^{*}$, the constrained generator can be implemented using a much smaller model, as its primary task is to rewrite the sketch $y^{*}$ with the help of constrained decoding, thus facilitating deployment on standard consumer-grade hardware.
@@
SGCD builds on the expectation that the constrained refined output $z^{*}$ should be at least as good as both $y^{*}$ (as $z^{*}$ respects the constraints) and $w^{*}$ (as $P_{\text{sk}}$ is a more powerful LLM than $P_{\text{cg}}$).
@@ 3 Experiments
     LLaMA-2-7B + CD 31.7± 2.7 36.8± 2.9 31.9± 2.5 36.6± 2.3 29.4± 2.0 31.2± 2.0
     LLaMA-2-13B + CD 31.4± 2.6 40.7± 3.0 33.1± 2.5 38.9± 2.4 32.6± 2.1 33.6± 2.1
Without logit access
     GPT4 43.2± 2.7 47.2±3.0 42.8± 2.6 46.8± 2.3 45.3± 2.2 45.8± 2.2
     GPT4 + SGCD 56.9± 3.2 (normal-$\uparrow$13.7) 47.4± 2.9 (normal-$\uparrow$2.1) 48.6± 2.8 (normal-$\uparrow$2.8) 53.2± 7.9 (normal-$\uparrow$3.5) 43.0± 2.3 (normal-$\downarrow$-2.3) 45.6± 2.2 (normal-$\downarrow$-0.2)
     GPT3.5 29.6± 2.7 27.0± 2.7 26.7± 2.5 25.7± 2.0 24.0± 1.9 24.6± 1.9
     GPT3.5 + SGCD 47.6± 3.3 (normal-$\uparrow$18.0) 37.9± 2.8 (normal-$\uparrow$10.9) 39.4± 2.8 (normal-$\uparrow$12.7) 44.2± 2.8 (normal-$\uparrow$18.5) 32.4± 2.1 (normal-$\uparrow$8.4) 35.3± 2.1 (normal-$\uparrow$10.7)
     Claude 37.6± 3.1 31.3± 2.8 32.5± 2.7 27.5± 2.0 27.6± 2.0 27.4± 2.0
     Claude + SGCD 49.4± 2.5 (normal-$\uparrow$11.8) 37.5± 2.8 (normal-$\uparrow$6.2) 40.3± 2.9 (normal-$\uparrow$7.8) 46.2± 2.5 (normal-$\uparrow$18.7) 35.8± 2.2 (normal-$\uparrow$8.2) 38.3± 2.2 (normal-$\uparrow$10.9)
     Claude-instant 25.3± 2.9 17.8± 2.2 19.9± 2.4 16.3± 1.7 15.9± 1.6 16.0± 1.6
     Claude-instant + SGCD 39.6± 3.3 (normal-$\uparrow$24.3) 29.6± 2.7 (normal-$\uparrow$11.8) 31.9± 2.5 (normal-$\uparrow$12.0) 37.8± 2.6 (normal-$\uparrow$21.5) 27.2± 2.1 (normal-$\uparrow$11.3) 30.0± 2.1 (normal-$\uparrow$14.0)
     PaLM-2-text 19.1± 1.9 23.0± 2.4 19.9± 2.0 28.1± 2.1 26.7± 2.0 27.1± 2.0
     PaLM-2-text + SGCD 43.5± 3.2 (normal-$\uparrow$24.4) 33.1± 2.7 (normal-$\uparrow$10.1) 35.2± 2.7 (normal-$\uparrow$15.3) 42.9± 2.6 (normal-$\uparrow$23.8) 31.9± 2.1 (normal-$\uparrow$8.9) 34.6± 2.1 (normal-$\uparrow$7.5)
     Average improvement normal-$\uparrow$16.4 normal-$\uparrow$7.9 normal-$\uparrow$10.7 normal-$\uparrow$15.4 normal-$\uparrow$5.8 normal-$\uparrow$8.1
Table 1: Results for closed information extraction, in terms of triplet-based precision, recall, and F1-score (micro-averaged, with bootstrapped 95% confidence intervals) on the Wiki-NRE and SynthIE-text datasets. The results compare the effectiveness of SGCD (blue rows) against two baselines: (1) few-shot-prompted unconstrained decoding with powerful blackbox LLMs (“without logit access”, white rows, Eq. 1) and (2) few-shot-prompted constrained decoding (“CD”) with open-source LLMs (“with logit access”, Eq. 3). Four demonstrations are used in few-shot prompting. LLaMA-7B serves as the constrained generator $P_{\text{cg}}$ for SGCD.
[…]
@@ 3.1 Closed information extraction
[…]
Constraints. We use the constraints in Appendix D.2, which restrict entities (1.5 million) and relations (857) to the Wikidata KG, and enforce the structural constraint that outputs must be formatted as sequences of entity–relation–entity triplets.
@@
Results. Table 1 illustrates that powerful blackbox LLMs, even without constrained decoding, already exceed the performance of open-source LLMs that employ constrained decoding. SGCD further enhances the performance of blackbox LLMs. Notably, precision shows a significant increase—by an average of 16.4% on Wiki-NRE and 15.4% on SynthIE-text—, demonstrating the efficacy of constrained decoding in producing outputs that are aligned with the KG. The gains in recall are more modest, with improvements of 7.9% and 5.8% on Wiki-NRE and SynthIE-text respectively, implying that constrained decoding primarily ensures alignment with the KG, rather than completeness. A detailed analysis of output grounding is provided in Appendix D.5.
Degeneration. Our method is built on the assumption that the sketcher will generate a high-quality sketch $y^{*}$, which the constrained generator will refine into $z^{*}$, meeting the specified constraints. Ideally, $z^{*}$ should match or surpass the quality of $y^{*}$, but in certain cases we observe a contrary outcome, which we term “degeneration”. This occurs when the constrained generator diverges from following the sketch $y^{*}$, opting to produce an alternative $z^{*}$ that, while meeting the constraints, is of inferior quality. Note (Table 1) that this issue leads to a (slight) decline in overall performance (in terms of recall) only with GPT-4 as the blackbox LLM, and only on the SynthIE-text dataset.
@@ 3.2 Constituency parsing
Penn Treebank
     Method B-precv B-recallv B-F1v Tag accuracyv Validity
@@
     LLaMA-2-7B + CD 8.2± 0.1 12.5± 0.2 9.9± 0.1 14.5± 0.1 92.0
     LLaMA-2-13B + CD 12.9± 0.1 20.2± 0.2 15.7± 0.1 16.6± 0.1 95.0
Without logit access
     GPT4 35.0± 0.1 30.6± 0.1 32.7± 0.1 47.0± 0.1 50.3
     GPT4 + SGCD 63.0± 0.1 (normal-$\uparrow$28.0) 57.4± 0.1 (normal-$\uparrow$26.8) 60.1± 0.1 (normal-$\uparrow$27.4) 89.5± 0.1 (normal-$\uparrow$42.5) 100 (normal-$\uparrow$49.7)
     GPT3.5 31.0± 0.1 24.6± 0.1 27.4± 0.1 47.1± 0.1 51.1
     GPT3.5 + SGCD 55.6± 0.1 (normal-$\uparrow$24.6) 44.9± 0.1 (normal-$\uparrow$20.3) 49.7± 0.1 (normal-$\uparrow$22.3) 87.3± 0.1 (normal-$\uparrow$40.2) 100 (normal-$\uparrow$48.9)
     Claude 22.4± 0.2 19.2± 0.1 20.7± 0.1 31.3± 0.1 33.2
     Claude + SGCD 47.4± 0.1 (normal-$\uparrow$25.0) 41.3± 0.1 (normal-$\uparrow$22.1) 44.1± 0.1 (normal-$\uparrow$23.4) 69.0± 0.2 (normal-$\uparrow$37.7) 99 (normal-$\uparrow$65.8)
     Claude-instant 23.6± 0.1 19.1± 0.1 21.1± 0.1 31.1± 0.1 34.6
     Claude-instant + SGCD 54.2± 0.1 (normal-$\uparrow$30.6) 44.7± 0.1 (normal-$\uparrow$25.6) 49.0± 0.1 (normal-$\uparrow$27.9) 77.2± 0.1 (normal-$\uparrow$46.1) 99.5 (normal-$\uparrow$64.9)
     PaLM-2-text 08.8± 0.5 06.7± 0.3 07.6± 0.3 13.2± 0.1 14.0
     PaLM-2-text + SGCD 60.9± 0.1 (normal-$\uparrow$52.1) 48.7± 0.1 (normal-$\uparrow$42.0) 54.1± 0.1 (normal-$\uparrow$46.5) 86.7± 0.1 (normal-$\uparrow$73.5) 100 (normal-$\uparrow$86)
     Average improvement normal-$\uparrow$32.1 normal-$\uparrow$27.4 normal-$\uparrow$29.4 normal-$\uparrow$48.0 normal-$\uparrow$63.1
Table 2: Results for constituency parsing, in terms of bracketing recall, precision, F1-score, and tag accuracy (with bootstrapped 95% confidence intervals), adjusted for validity (B-precv etc.), where each metric is scaled by the percentage of valid parse trees as per the EVALB tool. We compare the same methods as in constituency parsing (Table 1). Few-shot prompting uses three demonstrations.
@@
Task description. Constituency parsing involves breaking down a sentence into its syntactic components to form a parse tree that represents the sentence’s structure. For instance, the sentence “I saw a fox” corresponds to the parse tree [S [NP [PRP I]] [VP [VBD saw] [NP [DT a] [NN fox]]]]. For a visual representation of this tree, see Appendix E, Fig. 5. The instructions $I$ and $I_{\text{cg}}$ are listed in Appendix E.1.
@@
Constraints. We apply the constraints in Appendix E.2 to ensure that parse trees are complete, brackets are balanced, and labels are consistent.
[…]
Results. As indicated in Table 2, even advanced LLMs like GPT-4 face challenges in generating valid parse trees, particularly for longer sentences. LLaMA-2 models (7B, 13B), even using constrained decoding, exhibit much low performance, struggling predominantly with longer inputs Geng et al. (2023). However, with the aid of SGCD, the validity of output parse trees approaches 100% by construction (it may still be slightly less than 100% because the maximum output length is sometimes reached before the end of generation).
The validity adjustment also significantly enhances other metrics, reducing the penalty for invalid outputs. The most common errors in the unconstrained setting include unbalanced brackets and omissions. Appendix E, Fig. 7, provides a detailed example showing how SGCD effectively corrects these errors in draft parse trees.

### v3

@@ Abstract
Constrained decoding, a technique for enforcing constraints on language model outputs, offers a way to control text generation without retraining or architectural modifications. Its application is, however, typically restricted to models that give users access to next-token distributions (usually via softmax logits), which poses a limitation with blackbox large language models (LLMs). This paper introduces sketch-guided constrained decoding (SketchGCD), a novel approach to constrained decoding for blackbox LLMs, which operates without access to the logits of the blackbox LLM. SketchGCD utilizes a locally hosted auxiliary model to refine the output of an unconstrained blackbox LLM, effectively treating this initial output as a “sketch” for further elaboration. This approach is complementary to traditional logit-based techniques and enables the application of constrained decoding in settings where full model transparency is unavailable. We demonstrate the efficacy of SketchGCD through experiments in closed information extraction and constituency parsing, showing how it enhances the utility and flexibility of blackbox LLMs for complex NLP tasks. [^1] Code and data available at https://github.com/epfl-dlab/SketchGCD
Sketch-Guided Constrained Decoding for
Boosting Blackbox Large Language Models without Logit Access
Saibo Geng, Berkay Döner, Chris Wendler, Martin Josifoski, Robert West EPFL {saibo.geng, berkay.doner, chris.wendler, martin.josifoski, robert.west}@epfl.ch
@@ 1 Introduction
Figure 1: Overview of sketch-guided constrained decoding (SketchGCD). In the initial sketching phase, a blackbox LLM generates a preliminary “sketch” answer without applying any constraints. Then, in the constrained decoding phase, an auxiliary model, the constrained decoder, refines the sketch. The refined, final output respects the specified constraints by construction.
@@
[…]
@@ 2 Method
SketchGCD splits the constrained decoding task into two distinct phases: sketching and constrained decoding.
@@
Constrained decoding is done by a constrained decoder, a smaller-scale, locally hosted LLM $P_{\text{cg}}$. Given an instruction $I_{\text{cg}}$, a set of input–sketch–output demonstrations $D_{\text{cg}}=\{(x^{i},y^{i},z^{i})\}_{i=1}^{n}$, the original input $x$, and the sketch $y^{*}$, it refines $y^{*}$ into
@@
The sketcher’s output $y^{*}$ is typically of high quality, encapsulating the necessary information for the constrained decoder to produce the final sequence $z^{*}$ that adheres to the constraints $C$. Given the quality of $y^{*}$, the constrained decoder can be implemented using a much smaller model, as its primary task is to rewrite the sketch $y^{*}$ with the help of constrained decoding, thus facilitating deployment on standard consumer-grade hardware.
@@
SketchGCD builds on the expectation that the constrained refined output $z^{*}$ should be at least as good as both $y^{*}$ (as $z^{*}$ respects the constraints) and $w^{*}$ (as $P_{\text{sk}}$ is a more powerful LLM than $P_{\text{cg}}$).
@@ 3 Experiments
Without logit access
      GPT-4 42.4± 2.7 44.9±3.0 43.6± 2.6 46.1± 2.3 44.4± 2.2 45.2± 2.2
         + SketchGCD 7B 38.7± 3.2 ($\downarrow$3.7) 47.1± 2.9 ($\uparrow$2.2) 46.1± 2.8 ($\uparrow$2.5) 58.8± 7.9 ($\uparrow$12.7) 47.3± 2.3 ($\uparrow$2.9) 52.4± 2.2 ($\uparrow$7.2)
      GPT-3.5-Turbo 27.4± 2.0 27.4± 2.5 27.4± 2.5 24.6± 2.0 23.1± 1.9 23.8± 1.9
         + SketchGCD 7B 31.3± 3.3 ($\uparrow$3.9) 46.4± 2.8 ($\uparrow$18.7) 37.4± 2.8 ($\uparrow$10.0) 49.5± 2.8 ($\uparrow$24.9) 41.4± 2.1 ($\uparrow$18.3) 45.1± 2.1 ($\uparrow$21.3)
      Claude 34.1± 3.1 28.2± 2.8 30.8± 2.7 27.0± 2.0 26.7± 2.0 26.8± 2.0
         + SketchGCD 7B 30.4± 2.5 ($\downarrow$3.7) 40.6± 2.8 ($\uparrow$12.4) 34.8± 2.9 ($\uparrow$4.0) 51.4± 2.5 ($\uparrow$24.4) 36.3± 2.2 ($\uparrow$9.6) 42.5± 2.2 ($\uparrow$15.7)
      Claude-instant 24.5± 2.9 18.0± 2.2 20.8± 2.4 13.0± 1.7 15.2± 1.6 14.0± 1.6
         + SketchGCD 7B 44.9± 3.3 ($\uparrow$20.4) 31.1± 2.7 ($\uparrow$13.1) 36.7± 2.5 ($\uparrow$15.9) 44.9± 2.6 ($\uparrow$31.9) 31.1± 2.1 ($\uparrow$15.9) 36.7± 2.1 ($\uparrow$22.7)
@@
      LLaMA-2-7B 18.3± 2.4 14.0± 1.8 15.9± 1.2 12.0± 1.5 8.6± 1.1 10.0± 1.3
         + SketchGCD 7B 23.6± 2.7 ($\uparrow$5.3) 34.2± 2.9 ($\uparrow$20.2) 28.0± 2.4 ($\uparrow$12.1) 33.3± 2.5 ($\uparrow$21.3) 21.0± 2.0 ($\uparrow$12.4) 25.7± 2.1 ($\uparrow$15.7)
         + CD 33.6± 2.7 32.9± 2.9 32.8± 2.5 34.0± 2.3 25.9± 2.0 29.4± 2.0
      LLaMA-2-13B 22.6± 2.3 23.6± 2.4 23.1± 2.3 15.7± 1.6 12.7± 1.2 14.0± 1.5
         + SketchGCD 7B 28.8± 2.6 ($\uparrow$6.2) 44.2± 3.0 ($\uparrow$20.6) 34.9± 2.5 ($\uparrow$11.8) 36.1± 2.0 ($\uparrow$20.4) 25.1± 1.8 ($\uparrow$12.4) 29.6± 1.8 ($\uparrow$15.6)
         + CD 35.5± 2.6 39.1± 3.0 37.2± 2.5 39.7± 2.0 32.5± 1.8 35.7± 1.8
      LLaMA-2-70B 26.1± 2.7 24.5± 2.3 25.7± 2.4 32.6± 2.0 26.9± 1.8 29.4± 1.8
         + SketchGCD 7B 26.9± 2.7 ($\uparrow$0.8) 41.0± 2.6 ($\uparrow$16.5) 32.5± 2.1 ($\uparrow$6.8) 52.0± 2.0 ($\uparrow$19.4) 37.6± 1.8 ($\uparrow$10.7) 43.6± 2.0 ($\uparrow$14.2)
         + CD 39.9± 2.6 46.5± 2.6 42.3± 2.1 62.7± 2.0 50.3± 2.0 55.8± 2.0
Table 1: Results for closed information extraction, in terms of triplet-based precision, recall, and F1-score (micro-averaged, with bootstrapped 95% confidence intervals) on the Wiki-NRE and SynthIE-text datasets. The results compare the effectiveness of SketchGCD (blue rows) against two baselines: (1) few-shot-prompted unconstrained decoding with powerful blackbox LLMs (“without logit access”, white rows, Eq. 1) and (2) few-shot-prompted constrained decoding (“CD”) with open-source LLMs (“with logit access”, Eq. 3). Four demonstrations are used in few-shot prompting. LLaMA-7B serves as the constrained generator $P_{\text{cg}}$ for SketchGCD.
In our experimental setup, we evaluate the efficacy of SketchGCD by comparing it against two established baselines: (1) few-shot-prompted unconstrained decoding with powerful blackbox LLMs (Eq. 1) and (2) few-shot-prompted constrained decoding with open-source LLMs (Eq. 3). The SketchGCD method remains flexible and is agnostic to the exact implementation of constrained decoding. Here we adopt the grammar constrained decoding framework of Geng et al. (2023), but any other constraining method can be plugged in.
@@ 3.1 Closed information extraction
Task description. The goal of closed information triplet extraction (IE) is to extract a comprehensive set of facts from natural-language text. Formally, given a knowledge base represented by a knowledge graph (KG) containing a catalog of entities $\mathcal{E}$ and a catalog of relations $\mathcal{R}$, the goal is to extract the complete set $y_{\text{set}}\subset\mathcal{E}\times\mathcal{R}\times\mathcal{E}$ of fact triplets expressed in a given input text $x$. It is crucial that the entities and relations in these triplets be accurately grounded in the KG’s catalog. An example of this process can be seen in Fig. 1. The instructions $I$ and $I_{\text{cg}}$ for the sketcher and constrained decoder, respectively, are listed in Appendix D.1.
Constraints. We apply the constraints in Appendix D.2, which restrict entities (1.5 million) and relations (857) to the Wikidata KG, and enforce the structural constraint that outputs must be formatted as sequences of entity–relation–entity triplets.
@@
Results. We make the following observations based on Table 1: (1) The best blackbox LLMs (e.g., GPT-4) demonstrate strong performance even without constrained decoding, outperforming small open-source LLMs (LLaMA-2 7B/13B/33B) with constrained decoding. (2) Even without requiring access to logits, SketchGCD still manages to enhance the performance of LLMs significantly across all models of any size. (3) In case where logit access is available, constrained decoding is more effective than SketchGCD, as shown by the second half of the table. Given these observations, we conjecture that, if logits were accessible for blackbox LLMs, a further improvement in performance could be achieved with constrained decoding. However, without logit access, SketchGCD provides an effective alternative.
Impact of constrained decoder.
Wiki-NRE SynthIE-text
Prec Recall F1 Prec Recall F1
GPT-4 42.4 44.9 43.6 46.1 44.4 45.2
        + LLaMA-2-7B 38.7 57.1 46.1 58.9 47.3 52.4
        + LLaMA-2-13B 42.9 52.8 47.3 53.6 51.4 52.5
        + LLaMA-2-70B 35.2 54.0 42.6 58.1 53.1 55.5
Table 2: Impact of constrained decoder model (used in step 2 of SketchGCD) on closed information extraction. GPT-4 is used as the sketcher in all cases.
We investigate the impact of the constrained decoder on the performance of SketchGCD. As shown in Table 2, given GPT-4 as the sketcher, the choice of the constrained decoder can affect the performance of SketchGCD. Contrary to our expectations, larger constrained decoder models do not always lead to better performance. Our intuition is that step 2 of SketchGCD (constrained decoding) is relatively simple, and the additional capacity of larger constrained decoders does not necessarily provide an advantage.
Impact of beam size. Our experiments show that using beam search is critical for the performance of both SketchGCD and classical constrained decoding. As shown in Table 3, employing beam search (even with a minimal beam size of 2) significantly improves performance over greedy decoding. Larger beam sizes further enhance performance, allowing the model to explore a larger search space, but with a diminishing returns.
The following example illustrates the importance of beam search. Suppose we are doing closed information extraction on the sentence “Mona Lisa is housed in the Musée du Louvre in Paris.” Our entity catalog contains among other, the entities Louvre Museum and Musée d’Orsay. During unconstrained decoding, the model might generate the following output with highest probability: “[s] Mona Lisa [r] located in [o] Musée du Louvre. This output is invalid as the entity Musée du Louvre is not in the entity catalog and should be rendered as Louvre Museum instead.
With constrained decoding, the non-bold part of the output remains unaltered, as it satisfies the constraints. However, the bold suffix “du Louvre” is rejected by constrained decoding because Musée du Louvre is not in the entity catalog. The model will be forced to sample from the allowed entity catalog only, which can lead to “Musée d’Orsay” as the output. In this example, greedy constrained decoding was able to produce a valid yet incorrect output. On the contrary, had we used beam search, the model would have been able to consider both Musée du Louvre and Louvre Museum simultaneously, and would have been able to select the correct entity, Louvre Museum, for the output.
Wiki-NRE
LLaMA-2-7B + CD LLaMA-2-13B + CD
Prec Recall F1 Prec Recall F1
        1 beam 29.9 22.6 25.8 32.7 32.3 32.5
        2 beams 33.6 32.1 32.8 35.9 39.6 37.7
        4 beams 33.7 32.9 33.3 36.0 38.5 37.2
        8 beams 36.6 30.8 33.4 39.6 36.0 37.7
Table 3: Impact of beam size in beam search on closed information extraction during classical constrained decoding. “1 beam” is equivalent to greedy decoding.
@@ 3.2 Constituency parsing
       Method Bracket prec* Bracket recall* Bracket F1* Tag accuracy* Tag validity Tree validity
Without logit access
      GPT-4 76.6± 5.0 67.7± 4.5 71.9± 4.0 95.4± 0.9 93.6± 4.0 86.0± 4.0
          + SketchGCD 75.8± 2.4 ($\downarrow$0.8) 67.8± 2.4 ($\downarrow$0.1) 71.5± 2.4 ($\downarrow$0.4) 95.3± 0.8 ($\downarrow$0.1) 100± 0.0 ($\uparrow$6.4) 92.5± 4.0 ($\uparrow$6.5)
      GPT-3.5-Turbo 68.2± 0.7 55.5± 1.1 61.2± 0.6 93.1± 0.5 91.7± 2.4 76.9± 5.2
          + SketchGCD 68.7± 3.2 ($\uparrow$0.5) 56.6± 2.8 ($\uparrow$1.1) 62.1± 2.0 ($\uparrow$0.9) 92.6± 1.3 ($\downarrow$0.5) 100± 0.0 ($\uparrow$8.3) 81.5± 4.0 ($\uparrow$4.6)
      Claude 2.1 73.1± 3.3 63.1± 2.6 67.7± 2.5 94.5± 1.1 95.1± 2.5 62.6± 5.2
          + SketchGCD 71.6± 3.0 ($\downarrow$1.5) 62.9± 2.5 ($\downarrow$0.2) 66.9± 2.6 ($\downarrow$0.8) 93.4± 1.3 ($\downarrow$1.1) 100± 0.0 ($\uparrow$4.9) 68.7± 5.5 ($\uparrow$6.1)
      Claude-instant 1.2 71.3± 2.4 59.1± 1.4 64.7± 1.9 89.6± 1.6 91.7± 2.3 56.6± 5.2
          + SketchGCD 66.6± 3.3 ($\downarrow$4.7) 57.4± 3.1 ($\downarrow$1.7) 61.6± 3.3 ($\downarrow$3.1) 87.9± 2.5 ($\downarrow$1.7) 100± 0.0 ($\uparrow$8.3) 67.8± 3.7 ($\uparrow$11.2)
@@
       llama-2-7B 23.1± 4 10.4± 3 14.3± 4 14.9± 3 93.2± 3 32.1± 5
          + CD 28.5± 6 ($\uparrow$5.4) 16.5± 3 ($\uparrow$6.1) 20.9± 5 ($\uparrow$6.6) 13.8± 2 ($\downarrow$1.1) 100± 0 ($\uparrow$6.8) 35.1± 5 ($\uparrow$3.0)
       llama-2-13B 33.4± 7 22.4± 4 26.8± 5 29.3± 4 95.5± 2 38.5± 6
          + CD 33.3± 6 ($\downarrow$0.1) 21.8± 5 ($\downarrow$0.6) 26.3± 5 ($\downarrow$0.5) 34.0± 4 ($\uparrow$4.7) 100± 0 ($\uparrow$4.5) 43.4± 5 ($\uparrow$4.9)
       llama-2-70B 45.5± 6 37.7± 5 41.2± 5 $5$ 75.8± 5 40.4± 6
          + CD 39.8± 6 ($\downarrow$5.7) 35.6± 4 ($\downarrow$2.1) 37.6± 4 ($\downarrow$3.6) 53.8± 4 ($\downarrow$1.7) 100± 0 (($\uparrow$24.2)) 47.6± 5 ($\uparrow$7.2)
Table 4: Results for constituency parsing, in terms of bracketing precision, recall, F1-score, tag accuracy, tag validity, and parse tree validity (with bootstrapped 95% confidence intervals), on Penn Treebank test split. Only subset of samples whose ground-truth parse trees are shorter than 128 tokens (per LLaMA tokenizer) are considered (shortest 25% of the full dataset). Disclaimer: a weak method can have high precision by predicting very few valid parse trees (simple ones), and a strong method can have low precision by predicting more valid parse trees including complex ones (Deutsch et al., 2019). Four demonstrations are used in few-shot prompting. LLaMA-7B serves as the constrained generator $P_{\text{cg}}$ for SGCD. (* Considering only sentences with valid parse trees.)
@@
Task description. Constituency parsing involves breaking down a sentence into its syntactic components to form a parse tree that represents the sentence’s structure. For instance, the sentence “I saw a fox” corresponds to the parse tree [S [NP [PRP I]] [VP [VBD saw] [NP [DT a] [NN fox]]]]. For a visual representation of this tree, see Appendix E Fig. 5. The instructions $I$ and $I_{\text{cg}}$ are listed in Appendix E.1.
@@
Constraints. We apply the context-free grammar constraints in Appendix E.2 to ensure that brackets are balanced, and labels are consistent.
Dataset and evaluation metrics. Our evaluation uses the Penn Treebank test split. The parsing error rate of LLMs, regardless of size, is generally high, so we use only the shortest 25% of the samples for evaluation (up to 128 tokens according to the LLaMA tokenizer). We assess performance using bracketing recall and precision, as well as tag accuracy, as measured by the EVALB tool Sekine and Collins (2008). Since these metrics are only applicable to valid parse trees, and since models typically generate valid trees only for simpler inputs, one needs to be careful while interpreting the results, as weaker model may have better scores because they only generate a small fraction of valid parse trees (simpler ones) Deutsch et al. (2019).
Results. The results in Table 4 show that even advanced LLMs like GPT-4 struggle to generate valid parse trees, especially for longer sentences. The following observations can be made: (1) Both SketchGCD and classical constrained decoding significantly help the model generate more structurally valid parse trees. (2) The other metrics mostly remain unchanged or slightly drop, as a larger validity rate means more difficult examples are included in the evaluation. (3) The most common errors in the unconstrained setting are imbalanced brackets, invalid tags, and missing words, as shown in Table 5. (4) With SketchGCD, the error rate for imbalanced brackets and invalid tags is significantly reduced, while the error rate for missing words increases significantly.
Note that constrained decoding with a more sophisticated grammar, as described in Appendix E.2, can achieve 100% valid trees and 100% valid tags (see Table 9). However, as implementing such a grammar is non-trivial, we use a simpler context-free grammar here (see Appendix E.2) to mimic the real-world scenario where a simpler might be preferred over a perfect grammar.
Error type
Method InvalidTag Extra Imbal Missing
GPT-4 6.4% 0.4% 10.2% 2.3%
      + SketchGCD 0.0% 0.0% 2.6% 6.0%
GPT-3.5-Turbo 8.3% 2.6% 9.4% 2.3%
      + SketchGCD 0.0% 1.5% 1.9% 16.2%
Claude 2.1 4.9% 3.8% 3.4% 30.2%
      + SketchGCD 0.0% 3.0% 3.8% 29.8%
Table 5: Error analysis for constituency parsing on the Penn Treebank dataset. InvalidTag refers to model generating invalid tags, Extra to model adding extra words absent from input, Imbal to model generating imbalanced brackets, and Missing to model dropping words from input.